OpenStudy (anonymous):

the base of a cone has a radius of 6cm.The height of the cone is 15cm. What is the slant height of the cone? A.4.85 B. 16.16 C.19.21 D. 5.26

OpenStudy (kira_yamato):



hartnn (hartnn):

there is a pythagoras formula for this r^2+h^2=l^2 where r=base radius.,s=slant height, h= height of cone

OpenStudy (kira_yamato):

By Pythagorean Theorem, l^2 = 15^2 + 6^2

OpenStudy (anonymous):

Give a pictorial frame for the question

OpenStudy (kira_yamato):

@KKJ I think I kinda draw out the frame of the question in the drawing. Can you help me check if it's right @selena18 ? Arigato

OpenStudy (anonymous):

dont have a drawing for this problem

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 = 15^2 - 6^2 a ^2 = 189 a = 13.74

OpenStudy (anonymous):

is this right?

hartnn (hartnn):

nopes its 15^2+6^2 not minus.

hartnn (hartnn):

c=slant height = unknown

OpenStudy (anonymous):

15^2+6^2=261

OpenStudy (anonymous):

If l is the slant length, then l^2 = r^2 + h^2 where r and h are the radius and height of the cone resp.

hartnn (hartnn):

yup, now take square root of 261, what do u get ?

OpenStudy (anonymous):

its b

hartnn (hartnn):

yup

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

l^2 = 6^2 + 15^2 = 36 + 225 = 261 Hence , l = square root of 261 = 16.16 cm

hartnn (hartnn):

thank u so much for the medal, it was a special one from u, i became 99 :D

OpenStudy (anonymous):

The correct answer is B

OpenStudy (anonymous):

np =)